OpenStudy (anonymous):

A boat's crew rowed 10.5 miles downstream, with the current, in 1.5 hours. The return trip upstream , against the current , covered the same distance, but took 3.5 hours. Find the crew's rowing rate in still water and the rate of the current.

OpenStudy (anonymous):

Let x be the speed of the boat in still water and y the rate of the current. Then, \[Distance=Speed*Time\]When going downstream,\[10.5=(x+y)*1.5\]When going upstream,\[10.5=(x-y)*3.5\]Solving the two, we have x=5 and y=2

OpenStudy (anonymous):

Which is for the still water and which is for the the rate of the current?

OpenStudy (anonymous):

x is the speed of the crew in still water y is the rate of the current

OpenStudy (anonymous):

Okay thank you so much.